OpenStudy (mathmagician):

how much there will bee: 1/(1*2)+1/(2*3)+1/(3*4)+...+1/(99*100)? how much there will bee: 1/(1*2)+1/(2*3)+1/(3*4)+...+1/(99*100)? @Mathematics

OpenStudy (anonymous):

1-1/100=99/100

OpenStudy (mathmagician):

why is it so?

OpenStudy (anonymous):

1/1*2=(2-1)/1*2=1-1/2 similarly, 1/2*3=1/2-1/3 .... 1/99*100=1/99-1/100 if we add these, all the termas except 1 and 1/100 will be cancelled out

OpenStudy (cwrw238):

cddcdc

OpenStudy (mathmagician):

great, thanks :)

OpenStudy (anonymous):

sorry, terms, not termas :)

OpenStudy (jamesj):

Very nice.

OpenStudy (anonymous):

thanx